OpenStudy (anonymous):

how do i solve k^2+3k=108 ?

OpenStudy (anonymous):

k^2+3k+9/4=108+9/4 (k+3/2)^2=441/4 square root(k+3/2)^2= square root(441/4) k+3/2=21/2 k=9

OpenStudy (anonymous):

Wheww. Thanks Your'a life saver ( :

OpenStudy (anonymous):

hihih ur welcome..:)